OpenStudy (anonymous):

If g(x) = x2 − 5, find g(7)?

OpenStudy (anonymous):

7(2)-5=9

hero (hero):

g(x) = x^2-5 g(7) = 7^2-5 g(7) = 49-5 g(7) = 44

OpenStudy (anonymous):

Plug 7 into the x value

OpenStudy (anonymous):

Oh its squared?

hero (hero):

Yes

OpenStudy (anonymous):

If x is squared then hero has the correct answer and ignore mine